# Discrete Diffusion VLA: Bringing Discrete Diffusion to Action Decoding in Vision-Language-Action Policies

### v1

@@ 3.3 Unified VLA Architecture
Base architecture. We build on OpenVLA (Kim et al., 2024) as the representative backbone (Prismatic-7B VLM (Karamcheti et al., 2024) with SigLIP+DINOv2 (Zhai et al., 2023; Oquab et al., 2024b) visual encoders, a projector, and a Llama 2 language model (LM) backbone (Touvron et al., 2023a)). Unlike the original autoregressive action head, Discrete Diffusion VLA embeds discrete diffusion over actions inside the same transformer that processes vision and language, yielding a single, unified transformer as the VLA.
@@
Tokenization and action chunking. Following Brohan et al. (2022); Kim et al. (2024), we discretize each control dimension into a vocabulary of action tokens. As in OpenVLA, we adopt a 256-bin quantile-based scheme (1st–99th percentiles) per dimension to avoid outliers, and treat the gripper as a binary token. A single-timestep action therefore consists of $D_{\text{act}}{=}7$ tokens: 3 for translation, 3 for rotation, and 1 for gripper. For action chunking, we arrange tokens from $H$ future timesteps into a fixed layout, yielding a total of $L{=}H\!\times\!D_{\text{act}}$ action positions. We additionally append a special [MASK] token to the action vocabulary for discrete diffusion.
@@
Visual inputs comprise a mandatory third-person camera and optional left/right wrist cameras (Fig. 2). Each view is encoded by SigLIP and DINOv2 ViTs, whose features are projected into the LM embedding space; the language instruction is tokenized directly by the LM. Optional proprioceptive state (e.g., end effector positions and joint angles) is embedded via a small MLP and concatenated. The final input sequence is
[…]
where, during training, a randomly selected subset of the $L$ action positions is replaced by [MASK].
@@
Unified transformer and heads. All tokens pass through a single transformer, and we only predict logits at the action positions. For action tokens, we use a bidirectional attention mask which means no causal constraint and allow each action position to attend to all vision–language tokens. This design enables naturally full cross-modal fusion without bespoke adapters. Hidden states at action positions are projected to 256-way logits via a shared classification head. Vision and language follow standard VLM masking and heads if used.
@@
Compared with prior two action head designs, this unified backbone retains the language model’s representation power, scales seamlessly with model size, and allows parallel decoding. At inference time, our adaptive re-masking (detailed in Sec. 3.5) further refines uncertain tokens, combining the global-context strength of transformers with the iterative denoising spirit of diffusion.
@@ 3.4 Algorithmic Pipeline
Training pipeline. During training, for each minibatch, firstly we sample a mask ratio $\gamma\in(0,1]$ from a schedule (e.g., linear or cosine) that emulates diffusion time. Then we replace $\gamma L$ action positions with special token [MASK], and minimize the masked cross-entropy on those action indices following Eq. 6. We adopt hard-label supervision, representing ground truth as one-hot vectors at the masked indices. Vision and language tokens are left unchanged (standard LM masking/packing may be applied). This preserves the pretrained VLM capability since the objective is compatible with LM training, while aligning action learning with discrete diffusion via mask schedules. Notably, because the model predicts a fixed-length action chunk with $L{=}H\!\times\!D$ tokens, all sequences have uniform length and require no [EOS] padding. Joint optimization over the $L$ positions naturally trains the action chunk in a single pass.
Inference pipeline. At test time, we initialize all $L$ action positions as [MASK] and perform a small number of parallel refinement rounds. At each round $t$, our model predicts token posteriors for every currently masked position. We then draw a candidate token at each position by multinomial sampling from the predicted logits. Next, we set the mask ratio to $\gamma_{t}$ according to a preset schedule and use it to determine how many positions remain masked. We rank the masked positions by data-dependent scores (e.g., maximum confidence or confidence gap), commit the sampled tokens at the top $(1-\gamma_{t})$ fraction, and keep the remaining $\gamma_{t}$ fraction masked for the next round. This schedule makes the decoding order adaptive to each instance rather than fixed. The reverse-step conditionals follow the masking Bayesian formula in Eq. 5. A lightweight secondary re-masking mechanism further applies threshold and consistency checks to previously demasked tokens to prevent error propagation. Full criteria are detailed in Section 3.5.
@@ 3.5 Adaptive Decoding Mechanism and Secondary Re-Masking
Adaptive Decoding Mechanism. As illustrated above, the inference pipeline starts from a fully masked action chunk $\mathbf{a}_{1}=\mathrm{M}^{L}$ with mask ratio $\gamma_{1}{=}1$, and then performs $T$ refinement steps with a monotone schedule $\gamma_{t+1}<\gamma_{t}$. Here, we use cosine schedule following Chang et al. (2022b) as,
$\gamma_{t}=cos(\frac{\pi}{2}t),\ \ t\in[0,1)$ (7)
At step $t$, the model yields per-position posteriors $p_{\theta}(\mathbf{a}_{t-1}\mid\mathbf{a}_{t},\mathbf{c})$ instantiating the reverse conditionals in Eq. 5. We score each position $i$ using one of the adaptive metrics:
@@
with $k_{(1)},k_{(2)}$ the labels corresponding to the highest and second-highest probabilities. Let $m_{t,i}\!\in\!\{s_{t,i},\ g_{t,i}\}$. We keep the top $(1-\gamma_{t+1})L$ positions
$\mathcal{K}_{t}=\operatorname*{arg\,top}_{(1-\gamma_{t+1})L}\,m_{t,i},\vskip-2.0pt$ (10)
and update these positions’ tokens via tempered Gumbel sampling to encourage exploration:
@@
where $\bm{\varepsilon}$ has i.i.d. $\mathrm{Gumbel}(0,1)$ components (equivalently, Gumbel–Max), and $\tau_{t}$ is a temperature that decays with $\gamma_{t}$. Positions not in $\mathcal{K}_{t}$ are set to $\mathrm{M}$ (mask), and the process iterates until $\gamma_{T}{=}0$ or convergence. This instance-wise ranking makes the decoding order adaptive rather than fixed.
@@ 4.1 Benchmarks and Baselines
Benchmarks. We evaluate our Discrete Diffusion VLA on three different robot settings as shown in Fig. 3: (i) Franka Panda arm in LIBERO (Liu et al., 2023), using the four suites LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long (10 tasks per suite; 500 expert demos per suite). (ii) Google Robot in SimplerEnv–Fractal (Li et al., 2025), which reports Visual Matching and Variant Aggregation scores across diverse scene variations. (iii) WidowX Robot in SimplerEnv–Bridge, a real-to-sim evaluation aligned with BridgeData-V2 Walke et al. (2023) tasks. Policies receive only RGB images, that is one third-person view and one wrist view for LIBERO, and a single third-person view for SimplerEnv, together with a language instruction and optional end-effector positions; no depth, affordances, or other auxiliary information are used.
Baselines. We compare against representative policies spanning the two dominant paradigms, autoregressive (AR) token decoders and continuous diffusion / flow-matching policies, covering both models trained from scratch and models fine-tuned from large pretrained bases.
AR action decoders. RT-1-X / RT-2-X (O’Neill et al., 2024), OpenVLA (Kim et al., 2024), Octo-Small / Octo-Base (Ghosh et al., 2024), HPT (Wang et al., 2024), TraceVLA (Zheng et al., 2025), SpatialVLA (Qu et al., 2025), OpenVLA-OFT (Discrete) (Kim et al., 2025b) and $\pi_{0}$+FAST (Pertsch et al., 2025) represent AR-style generation of discrete action tokens with a unified VLM backbone.
Continuous diffusion / flow-matching. Diffusion Policy (Chi et al., 2023), MDT (Reuss et al., 2024), DiT Policy (Dita) (Hou et al., 2025), RoboVLM (Li et al., 2024b), $\pi_{0}$ (Black et al., 2024), OpenVLA-OFT (Cont.-Diffusion) (Kim et al., 2025b) and GR00T-N1 (Bjorck et al., 2025) instantiate denoising or flow-matching heads over continuous action trajectories. Seer is reported in both scratch and fine-tuned forms (Tian et al., 2025).
All baselines are evaluated on the three benchmarks using official metrics: LIBERO success rates; SimplerEnv–Fractal visual matching and variant aggregation success rates; and SimplerEnv–Bridge partial / full success rates. Unless otherwise noted, numbers are taken from the original papers or reproduced from open-source implementations under the same input modality described above.
4.2 Overall Performance Comparisons
Training details. We fine-tune Discrete Diffusion VLA on each benchmark from the same VLM backbone as OpenVLA (Prismatic–7B), following the respective official protocols. All input images are resized to $224\text{px}\times 224\text{px}$. For LIBERO, we train a separate policy per suite using the provided demonstrations, filtering unsuccessful episodes as in Kim et al. (2025b), and report success rates over the official test episodes. For SimplerEnv–Fractal and SimplerEnv–Bridge, we fine-tune on Fractal (Brohan et al., 2022) and BridgeData V2 (Walke et al., 2023), respectively. Across all settings, Discrete Diffusion VLA uses our unified transformer with discrete action tokens and a fixed action chunk. Chunk sizes are chosen to match the widely used settings for fair comparison: $8$ for LIBERO and SimplerEnv–Fractal but $5$ for SimplerEnv–Bridge. At inference, the adaptive decoding runs a small, fixed number of refinement rounds (12 by default) with a cosine mask schedule, which has been shown effective for discrete diffusion decoding (Chang et al., 2022b). Full hyperparameters and implementation details are provided in Appendix A.
@@ 4.2 Overall Performance Comparisons
OpenVLA-OFT (Cont-Diffusion) (Kim et al., 2025b) 96.9 98.1 95.5 91.1 95.4
@@
[…]
LIBERO results. Table 1 reports success rates (SR) on the four LIBERO suites. Discrete Diffusion VLA attains the best average SR of 96.3%, with per-suite scores of 97.2% (Spatial), 98.6% (Object), 97.4% (Goal), and 92.0% (Long). Our most comparable baseline is OpenVLA-OFT (Discrete), which uses the same action discretization as ours but decodes via parallel decoding rather than discrete diffusion. Discrete Diffusion VLA reaches 96.3% average SR vs 95.4% for OpenVLA-OFT (Discrete), a +0.9 point gain. These gains at matched tokenization indicate that discrete diffusion decoding provides a consistent advantage over parallel decoding. For reference, the pure AR baseline (OpenVLA) averages 76.5%, underscoring the benefit of moving beyond left-to-right decoding. Against methods trained from scratch, Discrete Diffusion VLA surpasses Diffusion Policy and MDT by +23.9 and +20.2 points on average, respectively. We observe consistent advantages across suites.
@@
Model Visual Matching Variant Aggregation Overall Average
@@
[…]
Google Robot results. As shown in Tab. 2, Discrete Diffusion VLA achieves the best Visual Matching average of 71.2%, clearly surpassing widely used baselines including $\pi_{0}$ (58.8%) and $\pi_{0}$+FAST (61.9%), and markedly outperforming OpenVLA-OFT (Discrete) (63.0% VM avg.). On Variant Aggregation, Discrete Diffusion VLA attains 56.9%, competitive with the top RT-2-X (64.3%) and $\pi_{0}$+FAST (59.0%). Aggregating both metrics, Discrete Diffusion VLA yields the highest overall average of 64.1%, reflecting strong robustness across appearance changes and task variants.
@@
Method Put Spoon on Towel Put Carrot on Plate Stack Green Block on Yellow #Overall
@@
Grasp Green Block
@@
RT-1-X (O’Neill et al., 2024) 16.7% 0.0% 20.8% 4.2% 8.3% 0.0% 8.3%
Octo-Base (Ghosh et al., 2024) 34.7% 12.5% 52.8% 8.3% 31.9% 0.0% 23.4%
Octo-Small (Ghosh et al., 2024) 77.8% 47.2% 27.8% 9.7% 40.3% 4.2% 34.5%
OpenVLA (Kim et al., 2024) 4.1% 0.0% 33.0% 0.0% 12.5% 0.0% 8.3%
RoboVLM (Li et al., 2024b) 54.2% 29.2% 25.0% 25.0% 45.8% 12.5% 32.0%
$\pi_{0}$ (Black et al., 2024) 45.8% 29.1% 25.0% 0.0% 50.0% 16.6% 27.8%
$\pi_{0}$-FAST (Pertsch et al., 2025) 62.5% 29.1% 58.5% 21.9% 54.0% 10.8% 39.5%
OpenVLA-OFT (Kim et al., 2025b) 50.0% 12.5% 41.7% 4.2% 70.8% 8.3% 31.3%
[…]
WidowX Robot results. On the WidowX evaluation in Tab. 3, Discrete Diffusion VLA attains the best overall average of 49.3%, outperforming diffusion / flow-matching policies ($\pi_{0}$: 39.5%, $\pi_{0}$+FAST: 39.5%) and exceeding AR-style baselines (e.g., Octo-Small at 34.5%). Per-task breakdown shows consistent gains in both grasp and success metrics (e.g., Put Spoon on Towel: 70.8% grasp / 37.5% success), indicating that discrete diffusion decoding improves reliability in these visually diverse manipulation settings.

### v2

@@ 3.3 Our Unified VLA Architecture
Base architecture. We build upon OpenVLA (Kim et al., 2024) architecture consisting of a Prismatic-7B VLM (Karamcheti et al., 2024) with SigLIP+DINOv2 (Zhai et al., 2023; Oquab et al., 2024b) visual encoders, a projector, and a Llama 2 language model (LM) backbone (Touvron et al., 2023a). Unlike the original autoregressive action head, Discrete Diffusion VLA modifies the causal-attention backbone into a bi-direction transformer that embeds discrete diffusion over actions inside the same VLM backbone, yielding a single, unified architecture.
@@
Tokenization and action chunking. Following RT series and OpenVLA (Brohan et al., 2022; Kim et al., 2024), we discretize each control dimension by a 256-bin quantile-based scheme (discretize only 1st–99th percentiles to avoid outliers), and treat the gripper as a seperate token between 0-1. A single-timestep action therefore consists of $D_{\text{act}}{=}7$ tokens: 3 for translation, 3 for rotation, and 1 for gripper. For action chunking, we arrange tokens from $H$ future timesteps into a fixed layout, yielding a total of $L{=}H\!\times\!D_{\text{act}}$ action positions. Discrete diffusion model excels at generating fixed length sequences. We additionally append a special [MASK] token to the action vocabulary.
@@
Visual inputs comprise a mandatory third-person image and two optional left and right wrist observations (Fig. 2). Each view is encoded by SigLIP and DINOv2 ViTs, whose features are projected into the Llama2 embedding space; the language instruction is tokenized directly by the Llama2. Optional proprioceptive states (i.e. end effector positions) are embedded via a small MLP and concatenated.
@@
Unified transformer and heads. All tokens (including vision, language and action) pass through the unified transformer. And we only predict logits at the action positions. For action tokens, we use a bidirectional attention mask which means no causal constraint and allow each action position to attend to all vision and language tokens. This design enables naturally full cross-modal fusion without bespoke adapters. Hidden states at action positions are projected to 256-way logit local vocabulary via a shared classification head. Vision and language tokens follow standard VLM masking.
@@
Compared to prior two action head designs, this unified backbone retains the language model’s representation ability, scales seamlessly with model size, and allows parallel decoding. At inference time, our adaptive re-masking (detailed in Sec. 3.5) further refines uncertain tokens, combining the global-context strength of transformers with the iterative denoising spirit of diffusion.
@@ 3.4 Algorithmic Pipeline
Training pipeline. During training, for each minibatch, firstly we sample a mask ratio $\gamma\in(0,1]$ from a schedule (e.g., linear or cosine) that emulates diffusion time. Then we replace $\gamma L$ action positions with special token [MASK], and minimize the masked cross-entropy on those action indices following Eq. 5. We adopt hard-label supervision, representing ground truth as one-hot vectors at the masked indices. Vision and language tokens are attended but ignored in loss. The objective is compatible with LM training which helps preserve the pretrained VLM capability, while aligning action generation with discrete diffusion via mask schedules. Because the model predicts a fixed-length action chunk with $L{=}H\!\times\!D$ tokens, all sequences have uniform length and require no [EOS] padding. Joint optimization over the $L$ positions naturally trains the action chunk in a single pass.
Inference pipeline. At test time, we initialize all $L$ action positions as [MASK] and perform a small number of parallel refinement rounds. At each round $t$, our model predicts token posteriors for every currently masked position. We then draw a candidate token at each position by multinomial sampling from the predicted logits. Next, we set the mask ratio to $\gamma_{t}$ according to a preset schedule and use it to determine how many positions remain masked. We rank the masked positions by data-dependent scores (e.g. maximum confidence or confidence gap), commit the sampled tokens at the top $(1-\gamma_{t})$ fraction, and keep the remaining $\gamma_{t}$ fraction masked for the next round. This schedule makes the decoding order adaptive to each instance rather than fixed. The reverse-step conditionals follow the masking Bayesian formula in Eq. 4. A lightweight secondary re-masking mechanism further applies threshold and consistency checks to previously demasked tokens to prevent error propagation. Full criteria are detailed in Section 3.5.
@@ 3.5 Adaptive Decoding Mechanism and Secondary Re-Masking
Adaptive Decoding Mechanism. As illustrated above, the inference pipeline starts from a fully masked action chunk $\mathbf{a}_{1}=\mathrm{M}^{L}$ with mask ratio $\gamma_{1}{=}1$, and then performs $T$ refinement steps with a monotone schedule $\gamma_{t+1}<\gamma_{t}$. Here, we use cosine schedule. At step $t$, the model yields per-position posteriors $p_{\theta}(\mathbf{a}_{t-1}\mid\mathbf{a}_{t},\mathbf{c})$ instantiating the reverse conditionals in Eq. 4. We score each position $i$ using one of the adaptive metrics:
@@
with $k_{(1)},k_{(2)}$ the labels corresponding to the highest and second-highest probabilities. Let $m_{t,i}\!\in\!\{s_{t,i},\ g_{t,i}\}$. We keep the top $(1-\gamma_{t+1})L$ positions $\mathcal{K}_{t}$ and update these positions’ tokens via tempered Gumbel sampling to encourage exploration:
@@
where $\bm{\varepsilon}$ has i.i.d. $\mathrm{Gumbel}(0,1)$ components (equivalently, Gumbel–Max), and $\tau_{t}$ is a temperature that decays with $\gamma_{t}$. Positions not in $\mathcal{K}_{t}$ are set to [MASK], and the process iterates until $\gamma_{T}{=}0$ or convergence. This instance-wise ranking makes the decoding order adaptive rather than fixed.
@@ 4.1 Benchmarks and Baselines
Benchmarks. We evaluate our Discrete Diffusion VLA on three different robot settings as shown in Fig. 3: (i) Franka Panda arm in LIBERO (Liu et al., 2023), using the four suites LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long (10 tasks per suite; 500 expert demos per suite). (ii) Google Robot in SimplerEnv–Fractal (Li et al., 2025), which reports Visual Matching and Variant Aggregation scores across diverse scene variations. (iii) WidowX Robot in SimplerEnv–Bridge, a real-to-sim evaluation aligned with BridgeData-V2 Walke et al. (2023) tasks. Policies receive only RGB images, that is one third-person view and one wrist view for LIBERO, and a single third-person view for SimplerEnv, together with a language instruction and optional end-effector positions. No depth, affordances, or other auxiliary information are used.
Baselines. We compare against representative policies spanning autoregressive (AR) token decoders, separate MLP action decoders and continuous diffusion / flow-matching action heads, covering both models trained from scratch and models fine-tuned from large pretrained bases.
Discretized action methods. RT-1-X / RT-2-X (O’Neill et al., 2024), OpenVLA (Kim et al., 2024), Octo-Small / Octo-Base (Ghosh et al., 2024), HPT (Wang et al., 2024), TraceVLA (Zheng et al., 2025), SpatialVLA (Qu et al., 2025), OpenVLA-OFT (Discrete) (Kim et al., 2025b) and $\pi_{0}$+FAST (Pertsch et al., 2025) represent AR-style or BERT-style generation of discrete action tokens with a unified VLM backbone or with a separate MLP decoder.
Continuous diffusion / flow-matching methods. Diffusion Policy (Chi et al., 2023), MDT (Reuss et al., 2024), DiT Policy (Dita) (Hou et al., 2025), RoboVLM (Li et al., 2024b), $\pi_{0}$ (Black et al., 2024), OpenVLA-OFT (Cont.-Diffusion) (Kim et al., 2025b) and GR00T-N1 (Bjorck et al., 2025) instantiate denoising or flow-matching heads over continuous action trajectories. Seer is reported in both scratch and fine-tuned forms (Tian et al., 2025).
All baselines are evaluated on at least one of the three benchmarks using official metrics: LIBERO success rates (SR); SimplerEnv–Fractal visual matching and variant aggregation SR; and SimplerEnv–Bridge partial and full SR. Unless otherwise noted, numbers are taken from the original papers or reproduced from open-source implementations under same input modality described above.
@@
OpenVLA-OFT (Cont-Diffusion) 96.9 98.1 95.5 91.1 95.4
@@
Discrete Diffusion VLA 97.2 98.6 97.4 92.0 96.3
4.2 Overall Performance Comparisons
Training details. We fine-tune Discrete Diffusion VLA on each benchmark from the same VLM backbone as OpenVLA (Prismatic–7B), following the respective official protocols. All input images are resized to $224\text{px}\times 224\text{px}$. For LIBERO, we train a separate policy per suite using the provided demonstrations, filtering unsuccessful episodes as in Kim et al. (2025b), and report success rates over the official test episodes. For SimplerEnv–Fractal and SimplerEnv–Bridge, we fine-tune on Fractal (Brohan et al., 2022) and BridgeData-V2 (Walke et al., 2023), respectively. Across all settings, Discrete Diffusion VLA uses our unified transformer with discrete action tokens and a fixed action chunk. Chunk sizes are chosen to match the widely used settings for fair comparison: $8$ for LIBERO and SimplerEnv–Fractal but $3$ for SimplerEnv–Bridge. At inference, the adaptive decoding runs a small, fixed number of refinement rounds (12 by default) with a cosine mask schedule, which has been shown effective for discrete diffusion decoding (Chang et al., 2022b).
LIBERO results. Table 1 reports success rates (SR) on four LIBERO suites. Discrete Diffusion VLA attains the best average SR of 96.3%, with per-suite scores of 97.2% (Spatial), 98.6% (Object), 97.4% (Goal), and 92.0% (Long). Our most comparable baseline is OpenVLA-OFT (Discrete), which uses the same action discretization as ours but decodes via parallel decoding rather than discrete diffusion. Discrete Diffusion VLA reaches 96.3% average SR vs 95.4% for OpenVLA-OFT (Discrete), a +0.9 point gain. These gains at matched tokenization indicate that discrete diffusion decoding provides a consistent advantage over parallel decoding. For reference, the pure AR baseline (OpenVLA) averages 76.5%, underscoring the benefit of moving beyond left-to-right decoding. Against methods trained from scratch, Discrete Diffusion VLA surpasses Diffusion Policy and MDT by +23.9 and +20.2 points on average, respectively. We observe consistent advantages across suites.
@@ 4.2 Overall Performance Comparisons
Model Visual Matching Variant Aggregation #Overall Average
@@
Discrete Diffusion VLA 85.4% 67.5% 60.6% 71.2% 82.5% 64.6% 23.6% 56.9% 64.1%
@@
Method Put Spoon on Towel Put Carrot on Plate Stack Green on Yellow Put Eggplant in Basket #Overall
@@
Grasp G Block
Success
Grasp Eggplant
@@
RT-1-X (O’Neill et al., 2024) 16.7% 0.0% 20.8% 4.2% 8.3% 0.0% 0.0% 0.0% 6.3%
Octo-Base (Ghosh et al., 2024) 34.7% 12.5% 52.8% 8.3% 31.9% 0.0% 66.7% 43.1% 31.3%
Octo-Small (Ghosh et al., 2024) 77.8% 47.2% 27.8% 9.7% 40.3% 4.2% 87.5% 56.9% 43.9%
OpenVLA (Kim et al., 2024) 4.1% 0.0% 33.0% 0.0% 12.5% 0.0% 8.3% 4.1% 7.8%
RoboVLM (Li et al., 2024b) 54.2% 29.2% 25.0% 25.0% 45.8% 12.5% 58.3% 58.3% 38.5%
$\pi_{0}$ (Black et al., 2024) 45.8% 29.1% 25.0% 0.0% 50.0% 16.7% 91.6% 62.5% 40.1%
$\pi_{0}$-FAST (Pertsch et al., 2025) 62.5% 29.1% 58.5% 21.9% 54.0% 10.8% 83.3% 66.6% 48.3%
OpenVLA-OFT (Kim et al., 2025b) 50.0% 12.5% 41.7% 4.2% 70.8% 8.3% 91.7% 37.5% 39.6%
GR00T-N1 Bjorck et al. (2025) 83.3% 62.5% 54.2% 45.8% 70.8% 16.7% 41.7% 20.8% 49.5%
Discrete Diffusion VLA 70.8% 29.2% 58.3% 29.2% 62.5% 20.8% 91.7% 70.8% 54.2%
Google Robot results. As shown in Tab. 2, Discrete Diffusion VLA achieves the best Visual Matching average of 71.2%, clearly surpassing widely used baselines including $\pi_{0}$ (58.8%) and $\pi_{0}$-FAST (61.9%), and markedly outperforming OpenVLA-OFT (63.0%). On Variant Aggregation, Discrete Diffusion VLA attains 56.9%, competitive with the top RT-2-X (64.3%) and $\pi_{0}$-FAST (59.0%). Aggregating both metrics, Discrete Diffusion VLA yields the highest overall average of 64.1%, reflecting strong robustness across tasks.
WidowX Robot results. On the WidowX evaluation in Tab. 3, Discrete Diffusion VLA attains the best overall average of 54.2%, outperforming diffusion / flow-matching policies ($\pi_{0}$: 40.1%) and exceeding AR-style and BERT-style baselines (e.g., Octo-Small at 43.9% and $\pi_{0}$-FAST: 48.3%). Per-task breakdown shows consistent gains in both grasp and success metrics (e.g., Put Eggplant in Basket: 91.7% grasp / 70.8% success), indicating that discrete diffusion decoding improves reliability in these visually diverse manipulation settings.